OpenStudy (anonymous):

Algebra 1, please help, will give medal. Which values of x and y will satisfy y = 5x and 12x + 9y = 60? A. x = 1, y = 19 B. x = 35, y = 20/19 C. x = 1/20, y = 1/90 D. x = 20/19, y = 100/19 E. x = 19, y = 50/19

OpenStudy (anonymous):

D. X=20/19, Y=100/19

Directrix (directrix):

@Anime_wolf Code of conduct: OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution http://openstudy.com/code-of-conduct